OpenStudy (anonymous):

Polygons PQRST and VWXYZ are similar. If PQ = 6 feet, QR = 11 feet, RS = 3 feet, ST = 7 feet, TP = 5 feet, and XY = 6 feet, then what is the length of WX?

OpenStudy (anonymous):

11/3 = x/6 6*11=3x 66=3x x=22 WX is 22 feet